OpenStudy (madmerc):

Describe how to transform the quantity of the sixth root of x to the fifth power, to the seventh powerinto an expression with a rational exponent. Make sure you respond with complete sentences.

zepdrix (zepdrix):

It's really hard to understand your question with the way it is written... Is it this? \(\large\rm \sqrt[6]{(x^5)^7}\) Or something else maybe? Hmm... I'm not sure. Err hmm, the quantity... so all of that quantity is to the seventh power. So something like this I think, \(\large\rm \left(\sqrt[6]{x^5}\right)^7\)